A postal clerk sold 40 stamps for $5.40. Some were 10-cent stamps and some were 15-cent stamps. How many 10-cent stamps were there?
Found 2 solutions by ikleyn, amalm06:
Answer by ikleyn(53937)

.
A postal clerk sold 40 stamps for $5.40. Some were 10-cent stamps and some were 15-cent stamps. How many 10-cent stamps were there?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Read the lesson What type of problems are these?, Problem 3, in this site.

This problem is very similar to yours.

Read its solution attentively.

Then solve you problem by substituting your data.

In this way you will learn how to solve problems like this, once and for all.


Lok also into associated lessons 

    Problem on animals at a farm 

    Problem on pills in containers     and 

    Solving coin problems without using equations 

in this site.


Answer by amalm06(224)

Let S1=10-cent stamps and S2=15-cent stamps
If 40 stamps cost $5.40, then 1 stamp cost 0.135
Using the method of alligation,
0.135-0.10=0.035
0.15-0.135=0.015
So that S1/S2=0.015/0.035
Then S1=(0.015/0.05)(40)=12 10-cent stamps (Answer)
S2=40-12=28 15-cent stamps (Answer)